OpenStudy (ayee_nya):

[Integers] Find the value of this integer expression: -2(3+-4+-8+7). Also, can someone please show the steps with distributive property and without. Thanks!

OpenStudy (eliesaab):

-4( 3-5+6)=-4(4)=-16 without distributive -4(3-5+6)= -12 +20-24=-16 with Distributive

OpenStudy (eliesaab):

Mimic the above to finish your problem

OpenStudy (ayee_nya):

Thanks!

OpenStudy (eliesaab):

YW